if sec theta = 13/12
what is sec theta/2?

Answer by lwsshak3(11628)

if sec x = 13/12
what is sec x/2?
**
work with more familiar cos function:
sec x = 13/12
cos x = 12/13 (reciprocal of sec x)
..
sec x/2=1/cos x/2
..
cos x/2=√[(1+cos x)/2]=√[(1+12/13)/2]=√[(25/13)/2]=√(25/26)=√25/√26
..
ans:
sec x/2=1/cos x/2=√26/√25